# Sparse Feature Coactivation Reveals Causal Semantic Modules in Large Language Models

### v2

@@ H.1 Individual Concept and Relation Ablation
Extending Table 1, we provide three comprehensive sets of examples for ablating individual concepts and relations: Table 9 (for country facts), Table 10 (for word translation), and Table 11 (for verb transformations). While the results largely mirror those highlighted in Table 1, we observe that when ablating the first letter transformation, the model still ranks tokens representing single letters highly, and many of those letters are other letters besides the first in the prompted verb (e.g., K, E, and A in the context of break). This suggests that the first letter component specifically captures first letters, and not other letters in prompted words; as such, ablating it still leaves a signal to the model to extract (non-first) letters.
Capital Currency Language
@@
Spanish French German
@@
Amor 41. " 5.9 Love 42. Amour 35. Le 15. Love 42. Liebe 66. Du 9.0 love 24.
amor 19. El 4.9 love 14. amour 26. La 6.4 I 11. Lie 4.8 " 6.6 Love 16.
Amo 7.0 La 4.4 I 13. A 5.7 " 6.1 love 10. Ich 4.4 Die 4.5 Liebe 11.
amo 5.1 Malo 3.2 " 5.3 J 3.8 le 3.7 " 5.8 " 3.4 das 4.2 I 9.6
" 2.9 la 2.5 LOVE 2.7 " 2.7 la 3.3 LOVE 2.5 Herz 1.7 Der 3.8 " 6.7
rojo 38. La 4.1 Red 29. Rouge 43. La 6.3 Red 38. Rot 51. " 5.2 red 30.
Rojo 34. " 3.5 red 16. rouge 31. " 6.1 red 20. rot 28. Gel 3.2 Red 18.
rojo 2.1 Bol 3.2 The 6.4 Rou 4.7 la 3.8 The 6.6 rote 5.8 Sch 2.5 The 6.1
Ro 2.0 la 2.8 " 6.4 ROU 2.6 Le 2.9 Rouge 5.8 Rote 3.5 1.9 " 5.8
Roja 1.9 2.7 4.1 Le 2.2 le 2.7 " 4.3 ro 2.0 die 1.5 2.9
@@
Past Tense Capitalize 1st Letter
@@ H.2 Individual Concept and Relation Steering
Extending Table 3, we provide three comprehensive sets of examples for steering individual concepts and relations: Table 12 (for country facts), Table 13 (for word translation), and Table 11 (for verb transformations). Observations from these extended examples largely resemble what we observed in Table 4. Related to what we observed in Table 11, we observe that steering toward capitalize is only partly successful at promoting the capitalized verb “Break,” while other letters in break besides the first letter (B) compete with it.
@@
_Abuja 87. _Yuan 15. _Mandarin 32. _Naira 75. _Beijing 98. _Chinese .35 _English 71. _Beijing 96. _Yuan 15.
_Nigeria 4.6 _yuan 12. _Chinese 27. _naira 8.3 _Be .75 _Mandarin .33 _Yoruba 5.6 _Be 1.5 _Ren 10.
_Lagos 3.9 _RMB 11. _English 25. _Nai 3.7 _BE .34 _English .28 _Ha 4.8 _Peking .83 _RMB 6.8
_- .58 _Ren 5.6 _Spanish 2.2 _Nigeria 3.3 _Peking .27 _Simplified .82 _Nigeria 4.4 Beijing .42 _yuan 6.3
_ .57 _China 5.4 _mandarin 2.1 _Nigerian 2.4 Beijing .25 _Spanish .79 _Igbo 3.4 _BE .30 _The 5.3
@@
_rojo 43. _A 11. _Liebe 65. _rouge 44. _amor 41. _Liebe 60. _rot 53. _amor 45. _A 11.
_Rojo 8.7 _l 9.3 _die 2.5 _Rouge 30. _el 9.5 _die 4.0 _Rot 33. _el 11. _l 9.3
_Sang 5.7 _amour 7.9 _zu 2.3 _Sang 2.5 _Amor 8.5 _„ 3.3 _rote 3.6 _" 7.4 _amour 7.9
_roja 2.8 _" 7.3 _„ 2.2 _Rou 1.8 _" 6.3 _lieben 1.8 _Rote 2.0 _a 3.2 _" 7.3
_rojo 1.8 _ 5.2 _Ich 1.7 _ 1.2 _A 2.8 _der 1.4 _Blut 1.2 _la 2.6 _ 5.2
@@
Past Tense, break Capitalize, break 1st Letter, break
$\cdot$, like Capitalize, $\cdot$ 1st Letter, $\cdot$ $\cdot$, like Past Tense, $\cdot$ 1st Letter, $\cdot$ $\cdot$, like Past Tense, $\cdot$ Capitalize, $\cdot$
_liked 43. _Past 17. _B 24. _Like 17. $\otimes$ 36. _break 19. _L 54. $\otimes$ 35. _" 6.8
_Liked 43. _Has 7.8 _( 8.2 _I 7.0 _surla 16. _B 18. _l 11. _surla 16. _K 6.4
_L 1.6 _Kick 4.4 _Past 5.8 _like 6.9 AddTagHelper 14. _b 7.6 _A 4.0 AddTagHelper 15. _Break 6.4
_Like 1.5 _ 4.2 _E 3.0 _LIKE 6.7 _These 7.0 _Break 6.1 _I 2.4 _These 7.2 _E 6.0
_- 1.1 _Have 4.0 _T 2.9 _The 6.0 _nahilalakip 5.7 _ 3.5 _" 2.1 _nahilalakip 6.3 _B 5.6
Table 14: Top 5 output tokens and likelihoods for prompts to get the past tense (columns 1–3), capitalized form (columns 4–6), and first letter (columns 7–9) of the word break, after ablating an in-prompt verb or transformation component (row 1 headings) and amplifying a target verb (red) or transformation component (row 2 headings). Highlighted tokens are or may begin correct answers for steered components. “s” refers to Old English long S.
@@ H.3 Composite Concept and Relation Steering
Extending Table 4, we provide two full sets of examples for steering concepts and relations in composition: Table 15 (for country facts) and Table 16 (for word translation). Observations from these extended examples largely resemble what we observed in Table 4.
@@
_Naira 34. _English 71. _Abuja 70. _English 93. _Abuja 49. _Naira 40.
_Nigeria 28. _French 5.1 _Lagos 26. _French 1.7 _Lagos 46. _Nigeria 18.
_naira 13. _Yoruba 4.1 _ 2.3 _Yoruba 1.5 _ 1.7 _Dollar 6.2
_N 4.1 _Spanish 3.8 _Nigeria .41 _Spanish 1.1 _Nigeria .78 _naira 4.7
_ 3.2 _Igbo 3.0 _… .30 _Igbo .55 _… .36 _ 3.7
@@
_rouge 38. _rot 31. _rojo 38. _rot 26. _rojo 45. _rouge 31.
_Rouge 5.2 _Rot 17. _el 9.1 _Rot 14. _el 9.8 _le 11.
_la 3.0 _rote 9.9 _roja 5.0 _rote 7.4 _roja 4.2 _la 4.9
_Sang 3.0 _die 3.7 _El 3.4 _die 4.6 _" 3.7 _ 4.5
_le 2.6 _der 1.8 _la 3.2 _„ 2.7 _la 2.6 _" 3.7
@@ Appendix I Steering Results by Concept
Table 2 summarized steering success rates by relation in each task. Here, Tables 17-19 break down concept steering success rate by concepts, while Tables 20-22 break down composite steering success rates by concept-relation pairs.
@@
Spanish 0.45 0.55 0.10 0.25 1.00 0.25 0.65 0.55 0.70 1.00 0.20 0.52
French 0.70 0.80 0.05 0.75 0.95 0.85 0.60 0.45 0.95 0.80 0.30 0.65
German 0.45 0.30 0.00 0.45 0.75 1.00 0.65 0.60 0.95 1.00 0.40 0.60
Average 0.53 0.55 0.05 0.48 0.90 0.70 0.63 0.53 0.87 0.93 0.30 0.59
@@
Capitalize 0.00 0.00 0.64 0.18 0.50 0.04 0.00 0.11 0.18
First Letter 0.93 0.36 0.93 0.86 1.00 1.00 1.00 1.00 0.88
Average 0.19 0.07 0.31 0.21 0.33 0.21 0.20 0.22 0.22
@@
Appendix J Supplementary Results on Component Organization
Extending Figure 4 from Section 3.3, we provide additional KL divergences by node and layer in Figure 13 (for translated words), Figure 14 (for target translation languages), Figure 15 (for verbs to transformer), and Figure 16 (for verb transformations).
@@
Appendix K Supplementary Results on Taxonomical Reasoning
@@
Appendix L Transcoder Analysis
@@
Appendix M License Information

### v3

@@ H.1 Individual Concept and Relation Ablation
Extending Table 1, we provide three comprehensive sets of examples for ablating individual concepts and relations: Table 12 (for country facts), Table 13 (for word translation), and Table 14 (for verb transformations). While the results largely mirror those highlighted in Table 1, we observe that when ablating the first letter transformation, the model still ranks tokens representing single letters highly, and many of those letters are other letters besides the first in the prompted verb (e.g., “K”, “E”, and “A” in the context of break). This suggests that the first letter component specifically captures first letters, and not other letters in prompted words; as such, ablating it still leaves a signal to the model to extract (non-first) letters. Further, it is possible that letter information is captured in a more general abstraction within the model which is activated by the first letter relation, but is not specific to it.
Capital, $\frac{\text{China}}{\text{Nigeria}}$ Currency, $\frac{\text{China}}{\text{Nigeria}}$ Language, $\frac{\text{China}}{\text{Nigeria}}$
@@
Spanish, $\frac{\text{love}}{\text{red}}$ French, $\frac{\text{love}}{\text{red}}$ German, $\frac{\text{love}}{\text{red}}$
@@
Amor 41. " 5.9 Love 42. Amour 35. Le 15. Love 42. Liebe 66. Du 9.0 love 16.
amor 19. El 4.9 love 14. amour 26. La 6.4 I 11. Lie 4.8 " 6.6 Liebe 13.
Amo 7.0 La 4.4 I 13. A 5.7 " 6.1 love 10. Ich 4.4 Die 4.4 Love 9.6
amo 5.1 Malo 3.2 " 5.3 J 3.8 le 3.7 " 5.8 " 3.4 das 4.1 I 8.7
" 2.9 la 2.5 LOVE 2.7 " 2.7 la 3.3 LOVE 2.5 Herz 1.7 Der 3.9 " 6.8
rojo 38. La 3.5 Red 27. Rouge 43. La 4.2 Red 36. Rot 51. " 4.6 red 30.
Rojo 34. 3.4 red 15. rouge 31. " 3.4 red 19. rot 28. Sch 2.8 Red 17.
rojo 2.1 " 3.2 " 8.1 Rou 4.7 3.2 The 5.9 rote 5.8 ge 2.5 rojo 6.1
Ro 2.0 El 2.8 The 5.1 ROU 2.6 Le 2.7 Rouge 5.6 Rote 3.5 2.2 Rojo 4.3
Roja 1.9 Bol 2.4 4.1 Le 2.2 En 2.6 " 5.5 ro 2.0 Gel 2.2 rouge 3.7
@@
Past Tense, $\frac{\text{break}}{\text{like}}$ Capitalize, $\frac{\text{break}}{\text{like}}$ 1st Letter, $\frac{\text{break}}{\text{like}}$
@@ H.2 Individual Concept and Relation Steering
Extending Table 3, we provide three comprehensive sets of examples for steering individual concepts and relations: Table 15 (for country facts), Table 16 (for word translation), and Table 14 (for verb transformations). Observations from these extended examples largely resemble what we observed in Table 3.
@@
_Abuja 87. _Yuan 13. _Mandarin 35. _Naira 75. _Beijing 97. _Chinese 38. _English 71. _Beijing 96. _Yuan 12.
_Nigeria 4.6 _yuan 11. _Chinese 28. _naira 8.3 _Be .94 _Mandarin 38. _Yoruba 5.6 _Be 2.0 _Ren 8.4
_Lagos 3.9 _RMB 11. _English 20. _Nai 3.7 _BE .57 _English 20. _Ha 4.8 _Peking .93 _China 6.1
_- .58 _China 7.0 _mandarin 2.4 _Nigeria 3.3 _Peking .33 _Simplified .92 _Nigeria 4.4 _BE .53 _Chinese 5.9
_ .57 _Ren 4.9 _Spanish 1.8 _Nigerian 2.4 Beijing .20 _Spanish .56 _Igbo 3.4 Beijing .37 _RMB 5.9
@@
_rojo 43. _amour 14. _Liebe 71. _rouge 44. _amor 41. _Liebe 68. _rot 53. _amor 50. _amour 14.
_Rojo 8.7 _A 11. _Ich 2.3 _Rouge 30. _Amor 13. _die 2.4 _Rot 33. _el 8.6 _" 9.7
_Sang 5.7 _" 6.7 _die 1.5 _Sang 2.5 _el 6.9 _„ 2.1 _rote 3.6 _" 7.4 _l 7.5
_roja 2.8 _l 6.5 _„ 1.5 _Rou 1.8 _" 6.1 _lieben 1.7 _Rote 2.0 _Amor 3.7 _A 6.7
_rojo 1.8 _Amour 5.3 _zu 1.4 _ 1.2 _A 3.1 _Ich 1.6 _Blut 1.2 _a 2.7 _ 6.5
@@
Past Tense, understand Capitalize, include 1st Letter, break
$\cdot$, break Cap., $\cdot$ 1st Lett., $\cdot$ $\cdot$, understand Past Tense, $\cdot$ 1st Lett., $\cdot$ $\cdot$, focus Past Tense, $\cdot$ Cap., $\cdot$
_broke 28. _Understand 6.7 _U 39. _Understand 14. $\otimes$ 35. _i 22. _F 26. $\otimes$ 30. _Break 21.
_Broke 21. _Know 6.4 _understood 9.3 _The 10. _surla 12. _I 21. _C 18. _surla 16.. _" 6.1
_BRO 8.5 _Past 5.9 _Understand 8.5 _I 6.3 AddTagHelper 13. _include 21. _O 15. AddTagHelper 13. _The 5.6
_Bre 6.3 _Answer 5.7 _u 3.6 _understand 4.3 _These 9.7 _Include 6.3 _f 4.4 _These 11. _“ 5.5
_brake 2.6 _ 4.5 _United 3.2 _ 4.1 _nahilalakip 6.9 _In 1.9 _o 3.3 _nahilalakip 8.6 _ 3.7
Table 17: Top 5 output tokens and likelihoods for prompts to get the past tense (columns 1–3), capitalized form (columns 4–6), and first letter (columns 7–9) of various verbs, after ablating an in-prompt verb or transformation component (row 1 headings) and amplifying a target verb (understand) or transformation component (row 2 headings). Highlighted tokens are or may begin correct answers for steered components. “s” refers to Old English long S.
@@ H.3 Composite Concept and Relation Steering
Extending Table 4, we provide two full sets of examples for steering concepts and relations in composition: Table 18 (for country facts) and Table 19 (for word translation). Observations from these extended examples largely resemble what we observed in Table 4. An unexpected exception is that the top next token when steering from the language of China to the capital of Nigeria is “Lagos” (another major city in Nigeria) rather than Abuja. We note that the full answer from the LLM contains the names of both cities, suggesting competition between these related tokens.
@@
_Nigeria 43. _English 75. _Abuja 70. _English 96. _Lagos 73. _Naira 48.
_Naira 30. _Yoruba 4.0 _Lagos 25. _Yoruba 1.1 _Abuja 23. _Nigeria 17.
_naira 8.9 _Igbo 3.6 _ 3.3 _French .74 _ 2.3 _Dollar 5.8
_N 3.0 _French 3.5 _Nigeria .34 _Spanish .60 _Nigeria .46 _ 4.5
_ 2.8 _Spanish 3.0 _… .32 _Igbo .54 _… .32 _naira 4.1
@@
_rouge 47. _rot 31. _rojo 43. _rot 28. _rojo 53. _rouge 45.
_Rouge 8.2 _Rot 26. _roja 6.5 _Rot 20. _el 7.2 _le 6.7
_Sang 3.2 _rote 11. _el 6.2 _rote 9.0 _roja 5.6 _Rouge 4.2
_sang 1.9 _die 2.1 _El 3.2 _die 2.9 _" 3.1 _ 3.7
_ 1.9 _Rote 1.6 _" 2.5 _roten 2.5 _El 2.2 _la 3.4
@@ Appendix I Steering Results by Concept
Table 2 summarized steering success rates by relation in each task. Here, Tables 20-22 break down concept steering success rate by concepts, while Tables 23-25 break down composite steering success rates by concept-relation pairs.
@@
Spanish 0.55 0.65 0.10 0.50 1.00 0.20 0.75 0.55 0.70 1.00 0.25 0.57
French 0.75 0.90 0.05 0.55 0.95 0.90 0.65 0.50 0.95 1.00 0.35 0.69
German 0.55 0.45 0.15 0.50 0.80 1.00 0.75 0.60 1.00 1.00 0.55 0.67
Average 0.62 0.67 0.10 0.52 0.92 0.70 0.72 0.55 0.88 1.00 0.38 0.64
@@
Capitalize 0.00 0.00 0.93 0.14 0.18 0.00 0.00 0.04 0.16
First Letter 1.00 0.00 0.89 0.93 0.71 1.00 1.00 0.75 0.79
Average 0.20 0.00 0.36 0.21 0.21 0.20 0.20 0.16 0.19
@@
Appendix J Steering Failure Analysis
In cases where steering fails, it is practically useful to understand the space of behaviors that may arise in LLMs. Figure 13 visualizes a categorization of steering failures quantified in Section 3.2.
Common failure cases.
In the country facts task, most failure cases were still partly successful, reinforcing the accuracy of our component selection in this task. Specifically, most incorrect counterfactual responses were correct answers for the target fact but not country (e.g., a capital city for the wrong country, often the prompted country). Additionally, many were at least relevant to target concepts or relations, e.g., names and non-capital cities of target countries. In word translation, most failures were similarly relevant to the target language and/or word. Usually, incorrect outputs were one or more non-target words in the target language. However, in rare cases the generated words were also nearly correct, e.g., steering from the French translation of love to run yielded “à pied” (meaning “on foot”). This suggests that for these two tasks, the concept and relation components we selected were generally fairly relevant to their corresponding contexts, even if they yielded some incorrect answers.
In verb transformation, however, most relation and composite steering failures invoked unintelligible or prompt-irrelevant model outputs. This again suggests that the components for this task are less reliable or specific than others, thus manipulating them can drastically impact general model performance. Notably, when steering from other relations to capitalization, the model often generated (incorrect) sequences of capitalized tokens, e.g., "The Word, The Word." This suggests that the capitalization component primarily represents stylization of generated outputs rather than word-specific application. In concept steering, which was more effective, most failures were partial successes, suggesting that the components for verbs are more viable, and still compose predictably with more complex transformation operations. Many cases yielded relevant responses for target relations and concepts, such as the target verb in the wrong form (e.g., not capitalized or in past tense), other letters besides the first letter of the target verb, and capitalized sequences of irrelevant tokens.
Nontrivial composition.
We additionally observed an interesting nontrivial composition of in-prompt and target concepts and relations in the word translation and verb transformation tasks.
In word translation, when steering from translating the word fish to other words across languages, the model often output translations of words related to water (e.g., ocean, sea, river, boat). This may suggest that the word fish is captured not only in our fish component, but also by other activated aquatic knowledge. When the fish component is ablated, this other knowledge may then integrate with target word knowledge in complex ways to yield responses that bridge the gap. For example, when steering from translating fish to here in both Spanish and German, the model generated Spanish/German words for sea, perhaps connecting fish to the location aspect of here. Meanwhile, when steering from translating fish to run in Spanish, the model generated a Spanish word for river (capable of “running”). While only anecdotal evidence, future work may consider further exploring how lexical knowledge is distributed across LLMs and synthesized during text generation.
In verb transformation, we again see nontrivial compositions of in-prompt and target relations, specifically when steering from other relations to capitalization. Here, steered outputs often consisted of capitalized tokens referring to the prompted relation, e.g., the model generated “The Past Tense Of The Word” when steering from past tense, “The Opposite of” when steering from antonym, and “The First Letter Of The” when steering from first letter. As mentioned in Section 3.2, this provides more evidence that the capitalization relation primarily represents the stylizing of generated text rather than the capability to manipulate it in a context-specific manner. Further, the intermingling of in-prompt and target information suggests that information about prompted relations is very widely distributed across many network layers and feature directions, even though the components for relations like past tense, antonym, and first letter are relatively large.
[…]
Appendix K Supplementary Results on Component Organization
Extending Figure 4 from Section 3.3, we provide additional KL divergences by node and layer in Figure 14 (for translated words), Figure 15 (for target translation languages), Figure 16 (for verbs to transform), and Figure 17 (for verb transformations).
@@
Appendix L Supplementary Results on Taxonomical Reasoning
@@
Appendix M Transcoder Analysis
@@
Appendix N License Information